-- RunaBriguglio - 26 Nov 2010

Optical test: T2 session

ps: for the complete description of this fancy day, see at the bottom of the page

act #454 has been removed from current

T2 setpoint:
Mirror command    
Bench command    
cold plate inlet    

tested S0, S1

test of int mat C0 and C1: before the flattenig shape S0 was applied

interaction matrix at T2

20101126_184001 0-250
20101126_184138 250-500
20101126_184310 501-652

'20101126_210700_noiseATTon' temperature inside the lab=9deg

20101126_235600 start testing at T1

tubone cooling off for testing

new int mat

20101127_003717 0-250
20101127_003838 250-500 dubbi
20101127_003948 501-652

20101127_004615 0-250
20101127_004849 250-500
0101127_005242 501-652

50% failures at the last steps

_______________________________________________

problems at startup:

error uploading DSP program to accel boards

errors on ADAM

new intmat calibrated at 7°
tracknum modes notes
083159 0-250 error curr
  251-500  
  501-653  
     
     
     

fanno tutte schifo

20101126_092249_flat650modes using the C2 command

and also 20101126_093127_flat650modes

sampled mode 650 for marco: 20101126_093458

stability test, loading shape 20101126_093127 (60nm)

start stability test 094516

Analizzato disturbo e dati presenti in

Analizzato disturbo e dati presenti in

'/towerdata/adsec_calib/disturbance/20101126_120525/diag_data.sav'

Scoperti attuatori 454ì-455 problematici.

Dopo una analisi tra posizioni e ffcmd integrato questa e' la mappa degli attuatori che non seguono bene:

  • Screenshot.png:
    Screenshot.png
Topic revision: r5 - 26 Nov 2010, MarcoXompero
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback